OpenStudy (anonymous):

A model ship is 28 in. by 65 in. Use the scale 1 in. : 1/2 ft to find the dimensions of the actual ship. (A)84 ft by 195 ft (B)14 ft by 65 ft (C)28 ft by 65 ft (D)14 ft by 32 and 1/2 ft

OpenStudy (anonymous):

You want to convert inches to feet given the scale 1 in per 1/2 ft.

OpenStudy (anonymous):

since you are going in to ft, you want in on the bottom and you want ft on top.

OpenStudy (anonymous):

\[ 28\, in \times \left(\frac{\frac{1}{2}\, ft}{1\, in}\right) = \quad ft \]

OpenStudy (anonymous):

the inches cancel and you are left with feet.

OpenStudy (anonymous):

When i did it i got B..but im not sure @wio

OpenStudy (anonymous):

You have to do the 65 in one too, do it the same way.

OpenStudy (anonymous):

i did n i got C @wio

OpenStudy (anonymous):

How could you get C? Can you show me your work?

OpenStudy (anonymous):

i did it on paper lol

OpenStudy (anonymous):

okay type it up a bit

OpenStudy (anonymous):

you can also draw it